OpenStudy (anonymous):

write an equation parallel to the line 1/2 y=3 through the point (-2,7). Please help

OpenStudy (anonymous):

y=7

OpenStudy (anonymous):

if (1/2)y=3, then solve for y=mx+b: y=3(2/1)=6 y=0x+6 therefore, the slope is 0 and any line with a slope of 0 passes through it. So a line that passes through (-2,7) would be y=7